A jet takes the same time to travel 2500 km with the wind as it does to travel 1900 km against the wind. If its speed relative to the air is 450 km/hr, what is the speed of the wind?
Answer by chrissi227(1)

Set up a system of equations and use the distance formula d = rt and solve for time.... t = d/r. The rate with the wind is just 450 + r and the rate against is 450 - r .... r is the rate of the wind.
Now you have two equations....
t = 2500/(450 + r)
t = 1900/(450 - r)
This leads you to....
2500/(450 + r) = 1900/(450 - r)
Cross multiply and distribute....
2500(450 - r) = 1900(450 + r)
1125000 - 2500r = 855000 + 1900r
Combine like terms....
4400 r = 270000
r = 61.36 km/h.
To 2 decimal places, the speed of the wind is 61.36 km/h.